# Second-Generation Planet Formation in Post-AGB Discs: Testing the Role of Gravitational Instability

### v1

@@ Abstract
Post-Asymptotic Giant Branch (post-AGB) binary stars are evolved systems that host circumbinary discs formed through mass loss during late-stage binary interaction. Their structural, morphological, kinematic, and chemical similarities to planet-forming discs have motivated the idea that these systems exhibit the characteristics of second-generation protoplanetary discs and may act as sites of “second-generation” planet formation. In this study, we assess whether the disc instability mechanism – a proposed pathway for giant planet formation in some protoplanetary discs—can operate in post-AGB discs. We test gravitational instability as a potential planet formation pathway, motivated by the short lifetimes ($10^{4-5}$years) of post-AGB discs, which would require any planet formation mechanism to operate efficiently on rapid timescales. Using the Toomre criterion under well-motivated assumptions for disc structure and size, mass, and thermal properties, we assess the conditions for gravitational instability. We first benchmark our analytical framework using well-studied protoplanetary disc systems (including HL Tauri, Elias 2-27, GQ Lupi) before applying the same analysis to observed post-AGB discs. We find that post-AGB discs are generally gravitationally stable at present, due primarily to their low masses. The disks were stable against collapse, even when accounting for potentially higher disk mass in the past. In contrast, several protoplanetary discs analysed in the same way show that they likely experienced gravitationally unstable phases early on. Finally, we revisit the planet formation scenario proposed for the post-common envelope system NN Serpentis first carried out by Schleicher and Dreizler and we show that gravitational instability could be feasible under specific, high disc mass assumptions, however we note that their proposed disc mass is unlikely. Overall, our results provide the first systematic theoretical assessment of gravitational instability in post-AGB discs, demonstrating that this mechanism is unlikely to dominate second-generation planet formation in these systems, and underscoring the need to explore alternative pathways—such as core accretion—in future studies.
@@
stars: AGB and post-AGB, planets and satellites: formation, protoplanetary disks, instabilities
@@ 1 Introduction
Understanding how and where planets form across different stages of stellar evolution remains one of the central questions in modern astrophysics. The Kepler and Gaia space telescopes have provided precise planet and stellar parameters for thousands of exoplanet systems 2010Borucki; 2018Thompson; 2016Gaia; 2020berger and high angular resolution observations with the Atacama Large Millimeter/ sub-millimeter Array (ALMA) have revealed direct signatures of ongoing planet formation in YSO discs, such as gaps, rings, cavities, and spirals - as seen in systems like PDS 70 Keppler_2019, HL Tauri Booth_2020, AB Aurigae Boccaletti_2020, Elias 2-27 Perez_2016, and HD 169142 Fedele_2017.
@@
Planets have also been detected or suspected around evolved stars. Notably, the first confirmed exoplanets were found orbiting the pulsar PSR B1257+12 — a system now known to host at least three planets 1992Wolszczan. The formation and/or survival pathway leading to planets existing in these extreme environments remains uncertain. For instance, 1993Podsiadlowski proposed that the disc from which pulsar planets formed originated through channels distinct from those forming YSO discs. Another example is the L2 Puppis system, an Asymptotic Giant Branch (AGB) star with an edge-on circumbinary disc and a candidate planetary companion – offering a glimpse into a potential future for our own solar system Kervella_2016. Similarly, NN Serpentis, a post-common envelope binary, exhibits eclipse timing variations interpreted as evidence for two circumbinary planets Beuermann_2010. These planets have been proposed to form from gas falling back following a common envelope phase 2014schleicher, as their low masses suggest that they may not have survived direct interaction with the RGB/AGB progenitor. However, alternative explanations — including magnetic activity cycles (the Applegate mechanism) and other dynamical effects — remain under debate Navarrete_2018; _zd_nmez_2023. Systems such as these point to the intriguing possibility of planet formation in previously unanticipated post-main sequence environments.
@@
Of specific interest to us are the discs around post-asymptotic giant branch (post-AGB) binaries. Post-AGB binaries are systems with a post-AGB star in the late stages of stellar evolution. The spectral energy distributions (SEDs) of the majority of post-AGB binary stars show a near-IR excess indicative of a circumbinary disc 2006deRuyter; 2014Kamath; Kamath_2015; Kluska_2022. The formation mechanisms of these discs are still poorly understood, but are generally presumed to have formed from late stage binary interactions, with more than 80 such discs now catalogued (kamath_2019; Kluska_2022; Andrych_2023, e.g.). The observational properties of these discs have been extensively studied and reveal striking similarities to those of protoplanetary discs around YSOs. For instance, interferometric maps with ALMA and observations of 12CO and 13CO lines Bujarrabal_2015; Gallardo_Cava_2023 have shown that many of these discs exhibit stable (quasi)-Keplerian dynamical properties 2006deRuyter; Gallardo_Cava_2021, with inner cavities in the dust component of the disk which is comparable to the dust sublimation radius in many of these systems Kluska_2019; Corporaal_2023. Furthermore, there is evidence of grain growth and dust grain evolution comparable to that seen in protoplanetary discs 2005deruyter; Scicluna_2020; Kluska_2022; Andrych_2023. Moreover, detailed chemical abundance studies of post-AGB binaries systems reveal a photospheric chemical depletion of refractory elements Kluska_2018, which has been interpreted as possibly being caused by a planet forming a dust-trapping gap in these discs. Photospheric depletions are stronger in post-AGB discs categorised as ”transition discs” mohorian2025tracingchemicaldepletionevolved, supporting the possibility that these discs harbour planets. Finally, the morphology of several ”transition” post-AGB discs includes features such as cavities (see Corporaal2023), with inner radii of up to 7.5 times the sublimation radius. These similarities have led to the hypothesis that post-AGB discs may also support planet formation.
@@
Planets around main sequence stars are presumed to form primarily via core accretion or gravitational instability (GI) drazkowska2023planetformationtheoryera. Core accretion is a bottom-up process involving the growth of small, grain-sized solids through coagulation and sticking during collisions and other mechanisms, eventually forming planetesimals and planetary embryos that can accrete additional material through gravitational interactions. The alternative theory, gravitational instability, is a top-down process in which massive regions of the disc undergo gravitational collapse, forming bound clumps that can contract and accrete material until planet-sized bodies emerge. This mechanism —similar in nature to the process of star formation— has been explored in seminal works such as 1997boss; 2002mayer. Gravitational instability is typically invoked to explain planet formation in the outer regions of discs, where cooling timescales and disc mass conditions may favour instability and fragmentation.
Gravitational instability can lead to planet formation on timescales comparable to the orbital period of the protoplanetary clump, making it a promising candidate for forming planets within short-lived post-AGB discs with lives of up to $10^{5}$ years Bujarrabal_2016. However, this mechanism also has limitations; most notably, it requires a sufficiently massive disc with the mass concentrated in a compact region for fragmentation to occur. Observationally, most post-AGB discs fall below this threshold. Nevertheless, 2014schleicher applied this mechanism to the NN Ser system, arguing that a relatively massive and compact disc – left behind after the CE phase – could have undergone gravitational fragmentation. Their models produced planet masses comparable to those observed in NN Ser, supporting the plausibility of second-generation planet formation via disc instability in select cases.
@@
This paper is structured as follows. In Section 2, we introduce the analytical framework used to assess the possibility of planet formation via gravitational instability. In Section 3, we apply this framework to a sample of young stellar object (YSO) discs that are known or suspected to host planets or exhibit signatures of gravitational instability, as an initial test of the theory. Section 4 presents the core focus of this work - an investigation of whether typical post-AGB discs can be gravitationally stable. In Section 5, we explore how the masses and sizes of present-day discs may have evolved over time and assess their early-life stability. Section 6 revisits the analysis by 2014schleicher of the NN Serpentis system, where we attempt to reproduce their calculations using more accurate assumptions about the disk’s temperature distributions.
@@ 2 Analytical Framework to Investigate Gravitational Instabilities
where $c_{\rm s}$ is the sound speed, $\Omega$ is the Keplerian angular frequency of the disc, $\Sigma$ is the disc’s surface density, $G$ is the gravitational constant , and $r$ is the radial distance from the center of the disc. While the classical Toomre criterion requires $Q\leq 1$ for instability against axisymmetric perturbations, simulations indicate that non-axisymmetric instabilities and spiral structures can emerge for $Q\leq 1.4$–1.7 Helled_2014.
The sound speed is related to the temperature, $T(r)$, by the following expression:
@@
where $k_{\rm B}$ and $m_{\rm p}$ are the Boltzmann constant and proton mass, respectively, and we have assumed the adiabatic index $\eta=1$  [^1] We denote the adiabatic index by $\eta$ to avoid confusion with the common use of $\gamma$ as a viscosity exponent in disc models. We adopt $\mu=2$ corresponding to molecular hydrogen gas.
@@
Chiang_1997, where $\alpha\approx 0.4R_{\rm star}/r$; $\alpha<<1$ is the grazing angle of the flux of stellar radiation incident on the disc, $R_{\rm star}$ is the photospheric radius of the star and $T_{\rm star}$ is the surface temperature of the star. This gives $T\propto r^{-0.75}$. This temperature profile is valid for optically thick discs in radiative equilibrium.
@@ 2.1 Cooling Efficiency
where $\eta=1.4$ for molecular Hydrogen-dominated gas, $\sigma$ is the Stefan–Boltzmann constant, and $f(\tau)=\tau+1/\tau$ is a function to quantify the efficiency of disc cooling, which takes the vertical optical depth to be $\tau=\kappa\Sigma/2$, where $\kappa$ is the opacity.
@@ 2.2 The Surface Density Profile
which is only valid for $n<2$. This type of function is used to fit the surface densities of both YSO discs 2023Speedie, and post-AGB discs Kluska_2022. For a fixed disc mass, decreasing $r_{\rm out}$ increases the overall surface density, effectively making the disc more compact (Figure 1).
@@ 2.2.1 Discs with Inner Cavities
Post-AGB discs have inner cavities not considered by Equations 10. Moreover, Equation 11 gives negative surface density values for power law indices $n\geq 2$ . We renormalise the surface density for discs with a finite inner radius, $r_{\rm in}$, resulting in the following:
@@
Figure 1: Power-law surface density functions with $n=1$ as a function of orbital distance for $M_{\rm disc}=0.15M_{\odot}$ for different values of $r_{\rm out}$, to demonstrate why the surface density is highly sensitive to $r_{\rm out}$.
@@ 2.2.2 Surface Density Profiles of Viscous Discs
While power-law models are commonly used to describe disc surface density profiles due to their simplicity, they are not physically motivated and rely on poorly defined parameters such as the outer radius, $r_{\rm out}$. Observationally, different diagnostics yield inconsistent estimates of $r_{\rm out}$ — for instance, optically thin dust continuum observations often indicate smaller disc sizes than optically thick molecular line tracers 2009Andrews.
@@ 3 Gravitational Instability in YSO Discs
Before applying our gravitational instability framework to the less explored environment of post-AGB discs, we first test its applicability using a set of well-studied YSO systems. Many of these systems show direct or indirect evidence of planet formation, and some exhibit structures–such as spirals, clumps, or large gaps–that have been associated with gravitational instabilities Cadman2021. By applying our analytical model to these systems, we aim to establish a benchmark for interpreting results in post-main-sequence discs.
@@
    GQ Lupi: a T Tauri star with a $10-40M_{\rm Jup}$ companion at $\sim 110AU$. While the current disc mass is relatively low ($\sim$3–4 $M_{\rm Jup}$), the wide separation of the companion suggests it may have formed via gravitational instability during an earlier, more massive phase of the disc Wu_2017. We note that the companion may lie in the brown dwarf regime, though its wide separation and possible disc origin make it relevant to GI-driven formation scenarios.
@@
In Table 1 we summarise the relevant properties of these systems used in our analysis. For AB Auriga and L1448 IRS3B systems, we assume a power-law surface density profile (for $\Sigma$, see Section 2.2), following 2023Speedie and Reynolds_2021. For HL Tauri and Elias 2-27, where stellar temperatures and radii are not available, we adopt plausible values consistent with their estimated stellar masses. We did not find an estimation for the inner radius of the disk of PDS-70 in the literature, hence we set its inner radius to $r_{\rm in}=0$.
In Figure 2, we plot the Toomre parameter $Q$ (see Section 2.1) as a function of disc radius for each system. While our focus here is on the Toomre criterion as a first-order diagnostic, we note that gravitational fragmentation also requires rapid cooling (see Section 2.1), which we do not assess in this section. We find that only two of the systems are Toomre-unstable at present (AB Aurigae with the disc mass reported in 2023Speedie, and L1448 IRS3B) indicating that gravitational instability is not generally active under current disc conditions for most of the systems, although some of them such as HL Tauri, GQ Lup, and Elias 2-27 are close to the unstable regime. We note that Toomre analyses have been carried out in the past for HL Tauri Booth_2020 and Elias 2-27 Perez_2016, and the resulting Toomre profiles in these studies match the profiles we found for them. This suggests that planet formation via gravitational instabilities may have occurred in these YSOs earlier in the disc’s evolution. We revisit this possibility in Section 5, where we explore whether these systems could have been gravitationally unstable at earlier stages.

### v2

# Second Generation Planet Formation in Post-AGB Discs: Testing the Role of Gravitational Instability

@@ Abstract
Post-asymptotic giant branch (post-AGB) binary stars are evolved systems that host circumbinary discs formed through mass loss during late stage binary interactions. Their structural, morphological, kinematic, and chemical similarities to planet-forming discs suggest that these systems may act as sites of “second generation” planet formation. In this study, we assess whether the disc instability mechanism – a proposed pathway for rapid, giant planet formation in some protoplanetary discs - can operate in post-AGB discs; motivated by their short lifetimes ($10^{4-5}$years). Using the Toomre criterion under well motivated assumptions for disc structure and size, mass, and thermal properties, we assess the conditions for gravitational instability. We first benchmark our analytical framework using well studied protoplanetary disc systems (including HL Tauri, Elias 2-27, GQ Lupi) before applying the same analysis to observed post-AGB discs. We find that post-AGB discs are generally gravitationally stable at present, due primarily to their low masses. Using viscous disk theory, we find that the discs were stable against collapse even in the past, when their masses were potentially higher. In contrast, several protoplanetary discs analysed in the same way show that they likely experienced gravitationally unstable phases early on. We also find that higher viscosity parameters ($\alpha\sim 10^{-2}$) are better aligned with expected post-AGB disc lifetimes. Finally, we revisit the planet formation scenario proposed for the post-common envelope system NN Ser, first carried out by Schleicher and Dreizler and we show that gravitational instability could be feasible under specific, high disc mass assumptions. Overall, our results provide the first systematic theoretical assessment of gravitational instability in post-AGB discs, demonstrating that this mechanism is unlikely to dominate second generation planet formation in these systems, and underscoring the need to explore alternative pathways - such as core accretion - in future studies.
@@
stars: AGB and post-AGB, planets and satellites: formation, protoplanetary discs, instabilities
@@ 1 Introduction
The Kepler and Gaia space telescopes have provided precise planet and stellar parameters for thousands of exoplanet systems 2010Borucki; 2018Thompson; 2016Gaia; 2020berger and high angular resolution observations with the Atacama Large Millimeter/ sub-millimeter Array (ALMA) have revealed direct signatures of ongoing planet formation in young stellar object (YSO) discs, such as gaps, rings, cavities, and spirals - as seen in systems like PDS 70 Keppler_2019, HL Tauri Booth_2020, AB Aurigae Boccaletti_2020, Elias 2-27 Perez_2016, and HD 169142 Fedele_2017.
@@
Planets have also been detected or suspected around evolved stars. Notably, the first confirmed exoplanets were found orbiting the pulsar PSR B1257+12 - a system now known to host at least three planets 1992Wolszczan. The formation and/or survival pathway leading to planets existing in these extreme environments remains uncertain. For instance, 1993Podsiadlowski proposed that the disc from which pulsar planets formed originated through channels distinct from those forming YSO discs. Another example is the L2 Puppis system, an asymptotic giant branch (AGB) star with a circumbinary disc and a candidate planetary companion - offering a glimpse into a potential future for our own solar system Kervella_2016. Similarly, NN Ser, a post-common envelope binary, exhibits eclipse timing variations interpreted as evidence for two circumbinary planets Beuermann_2010 (though alternative explanations for these variations such as magnetic activity cycles or other dynamical effects remain under debate 1992Applegate; Navarrete_2018; _zd_nmez_2023). These planets have been proposed to form from gas falling back following a common envelope phase 2014schleicher, as their low masses suggest that they may not have survived direct interaction with the RGB/AGB progenitor. . Systems such as these point to the intriguing possibility of planet formation in previously unanticipated post-main sequence environments.
@@
Of specific interest are the discs around post-asymptotic giant branch (post-AGB) binaries. Post-AGB binaries are systems with a post-AGB star in the late stages of stellar evolution. The spectral energy distributions (SEDs) of the majority of post-AGB binary stars show a near-IR excess indicative of a circumbinary disc 2006deRuyter; 2014Kamath; Kamath_2015; Kluska_2022. The formation mechanisms of these discs are still poorly understood, but they are presumed to have formed from late stage binary interactions, with more than 80 such discs now catalogued (kamath_2019; Kluska_2022; Andrych_2023, e.g.,). The observational properties of these discs have been extensively studied and reveal striking similarities to those of protoplanetary discs around YSOs. For instance, interferometric maps with ALMA and observations of 12CO and 13CO lines Bujarrabal_2015; Gallardo_Cava_2023 have shown that many of these discs exhibit stable (quasi)-Keplerian dynamical properties 2006deRuyter; Gallardo_Cava_2021. They also show inner cavities in the dust component of the disc that is comparable to the dust sublimation radius in many of these systems Kluska_2019; Corporaal_2023. Furthermore, there is evidence of grain growth and dust grain evolution comparable to that seen in protoplanetary discs 2005deruyter; Gielen_2011; Andrych_2023. Moreover, detailed chemical abundance studies of post-AGB stars in binary systems reveal a photospheric chemical depletion of refractory elements Kluska_2018, which has been interpreted as possibly being caused by a planet forming a dust-trapping gap in these discs Kluska_2022. Photospheric depletion is stronger in post-AGB discs categorised as ”transition discs” mohorian2025tracingchemicaldepletionevolved, supporting the possibility that these discs harbour planets. Finally, the morphology of several of these transition post-AGB discs includes features such as cavities Corporaal2023, with inner radii of up to 7.5 times the sublimation radius. These similarities have led to the hypothesis that post-AGB discs may also support planet formation.
@@
Planets around main sequence stars are presumed to form primarily via core accretion or gravitational instability (drazkowska2023planetformationtheoryera). Core accretion is a bottom-up process involving the growth of small, grain-sized solids through coagulation and sticking during collisions and other mechanisms, eventually forming planetesimals and planetary embryos that can accrete additional material through gravitational interactions. The alternative theory, gravitational instability, is a top-down process in which massive regions of the disc undergo gravitational collapse, forming bound clumps that can contract and accrete material until planet-sized bodies emerge. This mechanism - similar in nature to the process of star formation - has been explored in seminal works such as 1997boss; 2002mayer. Gravitational instability is typically invoked to explain planet formation in the outer regions of discs, where cooling timescales and disc mass conditions may favour instability and fragmentation.
Gravitational instability can lead to planet formation on timescales comparable to the orbital period of the protoplanetary clump, making it a promising candidate for forming planets within short-lived post-AGB discs with lives of up to $10^{5}$ years (Bujarrabal_2016, e.g.). However, this mechanism also has limitations; most notably, it requires a sufficiently massive disc with the mass concentrated in a compact region for fragmentation to occur. Nevertheless, 2014schleicher applied this mechanism to the NN Ser system, arguing that a relatively massive and compact disc – left behind after the CE phase – could have undergone gravitational fragmentation. Their model produced planet masses comparable to those observed in NN Ser, supporting the plausibility of second generation planet formation via disc instability in this case.
@@
This paper is structured as follows. In Section 2, we introduce the analytical framework used to assess the possibility of planet formation via gravitational instability. In Section 3, we apply this framework to a sample of YSO discs that are known or suspected to host planets or exhibit signatures of gravitational instability, as an initial test of the theory. Section 4 presents the core focus of this work - an investigation of whether typical post-AGB discs can be gravitationally unstable. In Section 5, we explore how the masses and sizes of present-day discs may have evolved over time and assess their early-life stability. Section 6 revisits the analysis by 2014schleicher of the NN Ser system, where we attempt to reproduce their calculations using more accurate assumptions about the disc’s temperature distributions.
@@ 2 Analytical Framework to Investigate Gravitational Instabilities
where $c_{\rm s}$ is the sound speed, $\Omega$ is the Keplerian angular frequency of the disc, $\Sigma$ is the disc’s surface density, $G$ is the gravitational constant , and $r$ is the radial distance from the center of the disc. While the classical Toomre criterion requires $Q\leq 1$ for instability against axisymmetric perturbations, simulations indicate that non-axisymmetric instabilities and spiral structures can emerge for $Q\leq 1.4$–1.7 Helled_2014. The sound speed is related to the temperature, $T(r)$, by the following expression:
@@
where $k_{\rm B}$ and $m_{\rm p}$ are the Boltzmann constant and proton mass, respectively, and we have assumed the adiabatic index $\eta=1$.  [^1] We denote the adiabatic index by $\eta$ to avoid confusion with the common use of $\gamma$ as a viscosity exponent in disc models. We adopt a mean molecular weight of $\mu=2$ corresponding to molecular hydrogen gas.
@@
Chiang_1997, where $\alpha\approx 0.4R_{\rm star}/r$ is the grazing angle of the flux of stellar radiation incident on the disc, and is typically much smaller than unity, $R_{\rm star}$ is the photospheric radius of the star and $T_{\rm star}$ is the surface temperature of the star. This gives $T\propto r^{-0.75}$. This temperature profile is valid for optically thick discs in radiative equilibrium.
@@ 2.1 Cooling Efficiency
where $\sigma$ is the Stefan–Boltzmann constant, and $f(\tau)=\tau+1/\tau$ is a function to quantify the efficiency of disc cooling, which takes the vertical optical depth to be $\tau=\kappa\Sigma/2$, where $\kappa$ is the opacity.
@@ 2.2 The Surface Density Profile
which is only valid for $n<2$. This type of function is used to fit the surface densities of both YSO discs 2023Speedie, and post-AGB discs Corporaal_2023. For a fixed disc mass, decreasing $r_{\rm out}$ increases the overall surface density, effectively making the disc more compact (Figure 1).
@@ 2.2.1 Discs with Inner Cavities
Post-AGB discs have inner cavities, evacuated regions, or holes centred on the central binary. These are not considered by Equation 10. Moreover, Equation 11 gives negative surface density values for power law indices $n\geq 2$ . To account for this if our disc has a cavity, we use Equation 11 but set $\Sigma(r)=0$ for $r<r_{\rm in}$, where $r_{\rm in}$ is the inner radius of the disc. The normalisation constant $\Sigma_{0}$ is consequently adjusted to account for a finite inner radius:
@@
Figure 1: Power-law surface density functions with $n=1$ as a function of orbital distance for $M_{\rm disc}=0.15~M_{\odot}$ for different values of $r_{\rm out}$, to demonstrate why the surface density is highly sensitive to $r_{\rm out}$.
@@ 2.2.2 Surface Density Profiles of Viscous Discs
While power-law models are commonly used to describe disc surface density profiles due to their simplicity, they are not physically motivated and rely on poorly defined parameters such as the outer radius, $r_{\rm out}$. Observationally, different diagnostics yield inconsistent estimates of $r_{\rm out}$ - for instance, optically thin dust continuum observations often indicate smaller disc sizes than optically thick molecular line tracers (2009Andrews, e.g.,).
@@ 3 Gravitational Instability in YSO Discs
Before applying our gravitational instability framework to the less explored environments of post-AGB discs, we first test its applicability using a set of well studied YSO systems. Many of these systems show direct or indirect features – such as spirals, clumps, or large gaps – that have been associated with gravitational instabilities Cadman2021. By applying our analytical model to these systems, we aim to establish a benchmark for interpreting results in post-main sequence discs.
@@
    GQ Lupi: a T Tauri star with a 104010-40 MJup companion at $\sim 110$ au. While the current disc mass is relatively low ($\sim$3–4 MJup), the wide separation of the companion suggests it may have formed via gravitational instability during an earlier, more massive phase of the disc Wu_2017. We note that the companion may lie in the brown dwarf regime, though its wide separation and possible disc origin make it relevant to gravitational instability-driven formation scenarios.
@@
In Table 1 we summarise the relevant properties of the YSOs used in our analysis, based on studies that have modelled the disc using observational data. For these systems, we have adopted the density profiles based on what was used in the respective studies, where some used the surface density profiles of viscous discs (Equation 13), while others used a power law surface density profile (Equation 12; see Table 1 and A for more details). For HL Tauri and Elias 2-27, where stellar temperatures and radii are not available, we adopt plausible values consistent with their estimated stellar masses. We did not find an estimation for the inner radius of the disc of PDS 70 in the literature, hence we set its inner radius to $r_{\rm in}=0$.
In Figure 2, we plot the Toomre parameter, $Q$ (see Section 2.1), as a function of disc radius for each system. While our focus here is on the Toomre criterion as a first-order diagnostic, we note that gravitational fragmentation also requires rapid cooling (see Section 2.1), which we do not assess in this section. We find that only two of the systems are Toomre-unstable at present (AB Aurigae with the disc mass reported in 2023Speedie, and L1448 IRS3B) indicating that gravitational instability is not generally active under current disc conditions for most of the systems, although some of them such as HL Tauri, GQ Lup, and Elias 2-27 are close to the unstable regime. We note that Toomre analyses have been carried out in the past for HL Tauri Booth_2020 and Elias 2-27 Perez_2016, and the resulting Toomre profiles in these studies match the profiles we found for them. As a further benchmark of our method, we compute the Toomre Q profiles of three protoplanetary discs with published Q profiles, as detailed in B.
Our findings suggest that planet formation via gravitational instabilities may have occurred in these YSOs earlier in the disc’s evolution. We revisit this possibility in Section 5, where we explore whether these systems could have been gravitationally unstable at earlier stages.